# Bi-Directional Deep Contextual Video Compression

### v1

@@ Abstract
Deep video compression has made remarkable process in recent years, with the majority of advancements concentrated on P-frame coding. Although efforts to enhance B-frame coding are ongoing, their compression performance is still far behind that of traditional bi-directional video codecs. In this paper, we introduce a bi-directional deep contextual video compression scheme tailored for B-frames, termed DCVC-B, to improve the compression performance of deep B-frame coding. Our scheme mainly has three key innovations. First, we develop a bi-directional motion difference context propagation method for effective motion difference coding, which significantly reduces the bit cost of bi-directional motions. Second, we propose a bi-directional contextual compression model and a corresponding bi-directional temporal entropy model, to make better use of the multi-scale temporal contexts. Third, we propose a hierarchical quality structure-based training strategy, leading to an effective bit allocation across large groups of pictures (GOP). Experimental results show that our DCVC-B achieves an average reduction of 26.6% in BD-Rate compared to the reference software for H.265/HEVC under random access conditions. Remarkably, it surpasses the performance of the H.266/VVC reference software on certain test datasets under the same configuration.
@@ I Introduction
The remainder of this paper is organized as follows. Section II gives a review of related work about deep video compression for P-frame and B-frame. Section III introduces a brief overview of our proposed DCVC-B scheme. Section IV describes our proposed methods in detail. Section V presents the experimental results and ablation studies. Section VI gives a conclusion of this paper.
@@ II-B Deep Video Compression for B-Frame
Figure 1: Overview of our proposed bi-directional deep contextual video compression scheme—DCVC-B. The motion estimation module estimates the bi-directional motion vectors ($v_{t\rightarrow f}$, $v_{t\rightarrow b}$) between the current frame $x_{t}$ and bi-directional reference frames ($\hat{x}_{f}$, $\hat{x}_{b}$) and also estimates the motion vector predictions ($v_{b\rightarrow f}$, $v_{f\rightarrow b}$) between ($\hat{x}_{f}$, $\hat{x}_{b}$). Then the motion vector differences (MVD) ($r_{t\rightarrow f}$, $r_{t\rightarrow b}$) between ($v_{t\rightarrow f}$, $v_{t\rightarrow b}$) and their predictions ($\frac{v_{b\rightarrow f}}{2}$, $\frac{v_{f\rightarrow b}}{2}$) are jointly compressed and decompressed by a motion encoder-decoder with our proposed bi-directional motion difference context propagation method. The reconstructed motion vectors ($\hat{v}_{t\rightarrow f}$, $\hat{v}_{t\rightarrow b}$) are used to perform bi-directional temporal context mining over the bi-directional reference features ($\hat{F}_{f}$, $\hat{F}_{b}$). The predicted bi-directional multi-scale temporal contexts $C_{f}^{0}$, $C_{f}^{1}$, $C_{f}^{2}$), ($C_{b}^{0}$, $C_{b}^{1}$, $C_{b}^{2}$) are fed into a contextual encoder-decoder to help compress and decompress the current frame $x_{t}$. Before obtaining the reconstructed frame $\hat{x}_{t}$, we regard an intermediate feature $\hat{F}_{t}$ of the contextual decoder as the propagated reference feature.
@@ IV-D Bi-directional Temporal Entropy Model
To further utilize the temporal predictions, we propose a bi-directional temporal entropy model, as shown in Fig. 7. The temporal entropy model learns a temporal prior from the bi-directional temporal contexts ($C_{t}^{f,2}$, $C_{t}^{b,2}$) and the bi-directional decoded latent representations ($\hat{y}_{t}^{f}$, $\hat{y}_{t}^{b}$). We first use two convolutional layers to obtain a feature $F_{t}^{c}$ with the same resolution and the same channel number as the current latent representation $\hat{y}_{t}$ from the bi-directional temporal contexts ($C_{t}^{f,2}$, $C_{t}^{b,2}$). Then we fuse $F_{t}^{c}$ with the bi-directional decoded latent representations ($\hat{y}_{t}^{f}$, $\hat{y}_{t}^{b}$) with a prior feature adaptor (PFA) and two depth blocks. The prior feature adaptor is also implemented by a depth block. For different types of latent representation propagation, we use different prior feature adaptors. When both the forward and backward reference frames are I-frames, as shown in Fig. 4(a), the $PFA_{0}$ is used. No decoded latent representation is fused.
$PF_{out}=PFA_{0}(F_{t}^{c}).$ (13)
When the forward reference frame is I-frames and the backward reference frame is B-frame, as shown in Fig. 4(b), the $PFA_{1}$ is used. The latent representation $\hat{y}_{t}^{b}$ of the backward reference frame is fused.
$PF_{out}=PFA_{1}(concat(F_{t}^{c},\hat{y}_{t}^{b})).$ (14)
When the forward reference frame is B-frames and the backward reference frame is I-frame, as shown in Fig. 4(c), the $PFA_{2}$ is used. The latent representation $\hat{y}_{t}^{f}$ of the forward reference frame is fused.
$PF_{out}=PFA_{1}(concat(F_{t}^{c},\hat{y}_{t}^{f})).$ (15)
When the forward and backward reference frames are both B-frames, as shown in Fig. 4(d), the $PFA_{3}$ is used. The latent representations ($\hat{y}_{t}^{f}$, $\hat{y}_{t}^{b}$) of the forward and backward reference frames are fused.
$PF_{out}=PFA_{3}(concat(F_{t}^{c},\hat{y}_{t}^{f},\hat{y}_{t}^{b})).$ (16)
@@ IV-E Hierarchical Quality Structure
To achieve a hierarchical quality structure for B-frame coding, traditional video codecs commonly assign different QPs (quantization parameters) to video frames of different temporal layers [55]. Recent deep P-frame coding schemes [6, 10] periodically increase the quality of video frames by assigning a quality coefficient that varies periodically over the frame index to the Lagrangian multiplier $\lambda$ in the rate-distortion (R-D) loss function. Inspired by traditional video codecs and deep P-frame coding schemes, we propose to assign different quality coefficients $w_{t}$ for the B-frames of different temporal layers. As shown in Fig. 2, we set $w_{t}$ of the B-frames in Layer 1, 2, 3, 4, and 5 to [1.4, 1.4, 0.7, 0.5, 0.5], respectively. When calculating the loss functions to train our model, we multiply $w_{t}$ on the $\lambda$. In addition, to make our model adapt to different qualities better, a quality adaptor implemented by one $1\times 1$ convolutional layer is inserted before the feature extractor module of the temporal context mining. For B-frames of different temporal layers, different quality adaptors are selected.
We combine the hierarchical quality structure-based training strategy with the step-by-step training strategy. The detailed training strategies are listed in Table. I. The definitions of the loss functions in different steps are described as follows.
@@ V-A1 Training Dataset
As listed in Table I, when the training frame is 3, 5, and 7, we train our model using 7-frame videos of the Vimeo-90k [56] dataset for short-sequence training. When the training frame is 17, we collect some sequences containing 33 frames from raw Vimeo videos. During training, the video frames are randomly cropped into 256$\times$256 patches.
@@ V-A3 Implementation Details
The DCVC-B scheme allows one model to support variable bitrates by introducing learnable quantization steps, as mentioned in Section IV-A and IV-C. During training, we set 4 base $\lambda$ values (85, 170, 380, 840) to control the rate-distortion trade-off. During testing, we can interpolate the quantization steps to achieve other bitrates. PyTorch is used to implement our model. AdamW [60] is used as the optimizer and the batch size is set to 8.
@@ V-A4 Test Configurations
In this paper, we focus on the random access (RA) scenario. For traditional video codecs, we choose the reference software of H.265/HEVC—HM-16.20 [61] and the reference software of H.266/VVC—VTM-13.2 [50] as our benchmarks. For HM, we use the default encoder_randomaccess_main_rext configuration (Intra period 32, GOP size 16). We denote HM under this configuration as HM-RA-GOP16. For VTM, we use the default encoder_randomaccess_vtm (Intra period 32, GOP size 16) and encoder_randomaccess_vtm_gop16.cfg (Intra period 32, GOP size 32) configurations. We denote VTM under these configurations as VTM-RA-GOP16 and VTM-RA-GOP32.
@@
For neural video coding models, we compare with existing state-of-the-art deep P-frame coding schemes, including DCVC-HEM [9], DCVC-DC [9], and DCVC-FM [12]. We do not compare with existing deep B-frame coding schemes since their best compression performance [14, 15] is still comparable with DCVC-HEM. The intra period and GOP size of our DCVC-B scheme and existing deep P-frame coding schemes are all set to 32. Different from [14] which compresses 97 frames of one video, we follow the existing deep P-frame coding schemes [10, 12] and compress 96 frames of one video to make a fair comparison. The structure of the first two GOPs is shown in Fig. 2, and the last one is the same as the VTM-RA-GOP32 except that it only has 2 reference frames (VTM has 4 reference frames).
@@ V-B1 Comparison Results
TABLE IV: Average encoding/decoding time for a 1080p frame (in seconds).
Schemes Enc Time Dec Time
@@
VTM-RA 1144.70 s 0.37 s
DCVC-HEM 0.75 s 0.26 s
DCVC-DC 0.82 s 0.64 s
DCVC-FM 0.74 s 0.53 s
Ours 1.19 s 0.99 s
V-B2 Encoding and Decoding Time
We list the encoding and decoding time for 1920$\times$1080 videos of different schemes in Table IV. When calculating the encoding and decoding time of deep video compression schemes, we follow the setting of [11, 6, 10, 9] and include the time for model inference, entropy modeling, entropy coding, and data transfer between CPU and GPU. We run deep video compression schemes on a NVIDIA 3090 GPU and run traditional video codecs on an Intel(R) Xeon(R) Gold 5118 CPU. The results show that the encoding and decoding time of our proposed DCVC-B scheme is increased compared with existing deep P-frame coding schemes. The main reason is the increased complexity of motion estimation on both sides of the encoder and decoder. We will try to reduce the model complexity in the future.
@@ V-B2 Encoding and Decoding Time
V-C Ablation Studies
V-C1 Effectiveness of Proposed Technologies
Figure 12: Visualization of the forward and backward temporal contexts of reference B-frames and non-reference B-frames.
@@ V-C2 Effectiveness of Bi-directional Motion Difference Context Propagation
V-C3 Effectiveness of Bi-directional Contextual Compression
V-C4 Effectiveness of Hierarchical Quality Structure
@@
VI Conclusion

### v2

@@ Abstract
Deep video compression has made impressive process in recent years, with the majority of advancements concentrated on P-frame coding. Although efforts to enhance B-frame coding are ongoing, their compression performance is still far behind that of traditional bi-directional video codecs. In this paper, we introduce a bi-directional deep contextual video compression scheme tailored for B-frames, termed DCVC-B, to improve the compression performance of deep B-frame coding. Our scheme mainly has three key innovations. First, we develop a bi-directional motion difference context propagation method for effective motion difference coding, which significantly reduces the bit cost of bi-directional motions. Second, we propose a bi-directional contextual compression model and a corresponding bi-directional temporal entropy model, to make better use of the multi-scale temporal contexts. Third, we propose a hierarchical quality structure-based training strategy, leading to an effective bit allocation across large groups of pictures (GOP). Experimental results show that our DCVC-B achieves an average reduction of 26.6% in BD-Rate compared to the reference software for H.265/HEVC under random access conditions. Remarkably, it surpasses the performance of the H.266/VVC reference software on certain test datasets under the same configuration. We anticipate our work can provide valuable insights and bring up deep B-frame coding to the next level.
@@ I Introduction
The remainder of this paper is organized as follows. Section II gives a review of related work about deep video compression for P-frame and B-frame. Section III introduces a brief overview of our proposed DCVC-B scheme. Section IV describes our proposed methods in detail. Section V presents the experimental results and ablation studies. Section VII gives a conclusion of this paper.
@@ II-B Deep Video Compression for B-Frame
Figure 1: Overview of our proposed bi-directional deep contextual video compression scheme—DCVC-B. The motion estimation module estimates the bi-directional motion vectors ($v_{t\rightarrow f}$, $v_{t\rightarrow b}$) between the current frame $x_{t}$ and bi-directional reference frames ($\hat{x}_{f}$, $\hat{x}_{b}$) and also estimates the motion vector predictions ($v_{b\rightarrow f}$, $v_{f\rightarrow b}$) between ($\hat{x}_{f}$, $\hat{x}_{b}$). Then the motion vector differences (MVD) ($r_{t\rightarrow f}$, $r_{t\rightarrow b}$) between ($v_{t\rightarrow f}$, $v_{t\rightarrow b}$) and their predictions ($\frac{v_{b\rightarrow f}}{2}$, $\frac{v_{f\rightarrow b}}{2}$) are jointly compressed and decompressed by a motion encoder-decoder with our proposed bi-directional motion difference context propagation method. The reconstructed motion vectors ($\hat{v}_{t\rightarrow f}$, $\hat{v}_{t\rightarrow b}$) are used to perform bi-directional temporal context mining over the bi-directional reference features ($\hat{F}_{f}$, $\hat{F}_{b}$). The predicted bi-directional multi-scale temporal contexts ($C_{f}^{0}$, $C_{f}^{1}$, $C_{f}^{2}$), ($C_{b}^{0}$, $C_{b}^{1}$, $C_{b}^{2}$) are fed into a contextual encoder-decoder to help compress and decompress the current frame $x_{t}$. Before obtaining the reconstructed frame $\hat{x}_{t}$, we regard an intermediate feature $\hat{F}_{t}$ of the contextual decoder as the propagated reference feature.
@@ IV-C Bi-directional Contextual Compression
Figure 8: Training reference structures for different numbers of training frames.
@@ IV-D Bi-directional Temporal Entropy Model
To further utilize the temporal predictions, we propose a bi-directional temporal entropy model, as shown in Fig. 7. The temporal entropy model learns a temporal prior from the smallest bi-directional temporal contexts ($C_{f}^{2}$, $C_{b}^{2}$) and the bi-directional decoded latent representations ($\hat{y}_{f}$, $\hat{y}_{b}$). We first use two convolutional layers to obtain a fused feature $C_{fb}^{2}$ with the same resolution and the same channel number as the current latent representation $\hat{y}_{t}$ from the bi-directional temporal contexts ($C_{f}^{2}$, $C_{b}^{2}$). Then we use a prior feature adaptor (PFA) and two depth blocks to fuse $C_{fb}^{2}$ with the bi-directional decoded latent representations ($\hat{y}_{f}$, $\hat{y}_{b}$). The prior feature adaptor is also implemented by a depth block. We use different prior feature adaptors for different types of latent representation propagation. When both the forward and backward reference frames are I-frames, as shown in Fig. 4(a), the $PFA_{0}$ is used. No decoded latent representation is fused.
$PF_{out}=PFA_{0}(C_{fb}^{2}).$ (13)
When the forward reference frame is I-frames and the backward reference frame is B-frame, as shown in Fig. 4(b), the $PFA_{1}$ is used. The latent representation $\hat{y}_{b}$ of the backward reference frame is fused.
$PF_{out}=PFA_{1}(concat(C_{fb}^{2},\hat{y}_{b})).$ (14)
When the forward reference frame is B-frames and the backward reference frame is I-frame, as shown in Fig. 4(c), the $PFA_{2}$ is used. The latent representation $\hat{y}_{f}$ of the forward reference frame is fused.
$PF_{out}=PFA_{1}(concat(C_{fb}^{2},\hat{y}_{f})).$ (15)
When the forward and backward reference frames are both B-frames, as shown in Fig. 4(d), the $PFA_{3}$ is used. The latent representations ($\hat{y}_{f}$, $\hat{y}_{b}$) of the forward and backward reference frames are fused.
$PF_{out}=PFA_{3}(concat(C_{fb}^{2},\hat{y}_{f},\hat{y}_{b})).$ (16)
@@ 17 All $L_{t}^{all}$ $65e-$ 6 2
Figure 9: Structure of the last GOP of our proposed DCVC-B scheme when compressing 96 frames. The structure is the same as VTM with the default encoder_randomaccess_vtm configuration (VTM-RA-GOP32) except that it has only 2 reference frames (VTM-RA-GOP32 has 4 reference frames).
@@ IV-E Hierarchical Quality Structure
To achieve a hierarchical quality structure for B-frame coding, traditional video codecs commonly assign different QPs (quantization parameters) to video frames of different temporal layers [55]. Recent deep P-frame coding schemes [6, 10] periodically increase the quality of video frames by assigning a quality coefficient that varies periodically over the frame index to the Lagrangian multiplier $\lambda$ in the rate-distortion (R-D) loss function. Inspired by traditional video codecs and deep P-frame coding schemes, we propose to assign different quality coefficients $w_{t}$ for the B-frames of different temporal layers. As shown in Fig. 2, we set $w_{t}$ of the B-frames in Layer 1, 2, 3, 4, and 5 to [1.4, 1.4, 0.7, 0.5, 0.5], respectively. When calculating the loss functions to train our model, we multiply $w_{t}$ on the $\lambda$. In addition, to make our model adapt to different qualities better, a quality adaptor implemented by one $1\times 1$ convolutional layer is inserted before the feature extractor module of the temporal context mining. For B-frames of different temporal layers, different quality adaptors are selected. Note that the quality adaptors are included for training from the beginning.
We combine the hierarchical quality structure-based training strategy with the step-by-step training strategy. The detailed training strategies are listed in Table. I. For different numbers of training frames, the training reference structures are illustrated in Fig. 8. The definitions of the loss functions in different steps are described as follows.
@@
  • †Due to the MS-SSIM model of DCVC-FM is not released, when the reconstruction quality is measured by MS-SSIM, we use the PSNR model of DCVC-FM to supplement the result, although this is unfair.
@@ V-A1 Training Dataset
As listed in Table I, when the training frame is 3, 5, and 7, we train our model using 7-frame videos of the Vimeo-90k [56] dataset for short-sequence training. When the training frame is 17, we use 9000 33-frame video clips collected from raw Vimeo videos. During training, the video frames are randomly cropped into 256$\times$256 patches.
@@ V-A3 Implementation Details
The DCVC-B scheme allows one model to support variable bitrates by introducing learnable quantization steps, as mentioned in Section IV-A and IV-C. During training, when the reconstruction quality is measured by PSNR, we set 4 base $\lambda$ values (85, 170, 380, 840) to control the rate-distortion trade-off and set the distortion metrics ($D_{t}^{y}$, $D_{t}^{m}$) to Mean Squared Error (MSE). When the reconstruction quality is measured by MS-SSIM, we set the base $\lambda$ values to ($\frac{85}{17}$, $\frac{170}{17}$, $\frac{380}{17}$, $\frac{840}{17}$) and set the distortion metrics ($D_{t}^{y}$, $D_{t}^{m}$) to 1–MS-SSIM. Since existing deep video coding schemes have outperformed traditional video codecs by a large margin regarding MS-SSIM, we only fine-tune our MS-SSIM model for 2 epochs based on the PSNR model. During testing, we can interpolate the quantization steps to achieve other bitrates. PyTorch is used to implement our model. AdamW [60] is used as the optimizer and the batch size is set to 8.
@@ V-A4 Test Configurations
In this paper, we focus on the random access (RA) scenario. For traditional video codecs, we choose the reference software of H.265/HEVC—HM-16.20 [61] and the reference software of H.266/VVC—VTM-13.2 [50] as our benchmarks. For HM, we use the default encoder_randomaccess_main_rext configuration (Intra period 32, GOP size 16). We denote HM under this configuration as HM-RA-GOP16. For VTM, we use the default encoder_randomaccess_vtm (Intra period 32, GOP size 32) and encoder_randomaccess_vtm_gop16.cfg (Intra period 32, GOP size 16) configurations. We denote VTM under these configurations as VTM-RA-GOP16 and VTM-RA-GOP32.
@@
For neural video coding models, we compare with existing state-of-the-art deep P-frame coding schemes, including DCVC-HEM [9], DCVC-DC [9], and DCVC-FM [12]. We do not compare with existing deep B-frame coding schemes since their best compression performance [14, 15] is still comparable with DCVC-HEM. The intra period and GOP size of our DCVC-B scheme and existing deep P-frame coding schemes are all set to 32. Different from [14] which compresses 97 frames of one video, we follow the existing deep P-frame coding schemes [10, 12] and compress 96 frames of one video to make a fair comparison. The structure of the first two GOPs is shown in Fig. 2. The structure of the last GOP is shown in Fig. 9, which is the same as the VTM-RA-GOP32 except that it has only 2 reference frames (VTM-RA-GOP32 has 4 reference frames). We use the same I-frame codec as DCVC-DC.
@@ V-B1 Comparison Results
TABLE IV: Runtime and computational complexity comparison for a 1080p video frame.
Schemes Enc Time Dec Time MACs/pixel Model Size
@@
VTM-RA 1144.70 s 0.37 s —-
DCVC-HEM 0.75 s 0.26 s 1791.64K 17.52M
DCVC-DC 0.82 s 0.64 s 1397.90K 18.45M
DCVC-FM 0.74 s 0.53 s 1180.77K 17.015M
Ours 1.19 s 0.99 s 3004.52K 22.28M
V-B2 Runtime and Computational Complexity Comparison
We list the encoding and decoding time for 1920$\times$1080 videos of different schemes in Table IV. When calculating the encoding and decoding time of deep video compression schemes, we follow the setting of [11, 6, 10, 9] and include the time for model inference, entropy modeling, entropy coding, and data transfer between CPU and GPU. We also compare the computational complexities and model sizes of deep video compression schemes in Table IV. We run deep video compression schemes on an NVIDIA 3090 GPU and run traditional video codecs on an Intel(R) Xeon(R) Gold 5118 CPU. The comparison results show that the complexity of our proposed DCVC-B scheme is increased compared with existing deep P-frame coding schemes. The main reason is the increased complexity of bi-directional motion estimation on both the encoder and decoder. We will try our best to reduce the complexity in the future.
V-C Ablation Studies
V-C1 Effectiveness of Proposed Technologies
@@ V-C2 Effectiveness of Bi-directional Motion Difference Context Propagation
Figure 14: Visualization of the forward and backward temporal contexts of different B-frames.
@@
V-C3 Effectiveness of Bi-directional Contextual Compression
V-C4 Effectiveness of Hierarchical Quality Structure
Figure 16: Frame quality and bitrate comparison between our proposed DCVC-B and DCVC-DC when the GOP size is set to 64.
TABLE VII: BD-rate (%) comparison of DCVC-DC and our DCVC-B over each testing sequence. The anchor HM-RA-GOP16.
Datasets Sequences DCVC-DC DCVC-B Datasets Sequences DCVC-DC DCVC-B
HEVC Class B Cactus –31.0 –37.2 MCL-JCV videoSRC01 –17.4 –33.0
ParkScene –9.9 –30.8 videoSRC02 –16.0 0.4
Kimono1 –29.2 –29.7 videoSRC03 –23.4 –37.4
BQTerrace –12.2 –20.5 videoSRC04 –19.5 –13.4
BasketballDrive –26.3 –11.8 videoSRC05 –30.0 –11.3
HEVC Class C BQMall –20.3 –14.8 videoSRC06 –28.8 56.1
BasketballDrill –25.1 –26.9 videoSRC07 –29.9 –29.9
RaceHorses –22.0 –0.8 videoSRC08 –31.7 –20.8
PartyScene –15.4 –23.7 videoSRC09 –26.6 –40.1
HEVC Class D BasketballPass –36.4 –38.4 videoSRC10 -13.6 4.6
BQSquare –22.0 –42.4 videoSRC11 –35.5 –7.6
BlowingBubbles –24.5 –37.3 videoSRC12 –1.3 –31.3
RaceHorses –42.4 –34.9 videoSRC13 19.6 –18.0
HEVC Class E KristenAndSara –22.8 –43.1 videoSRC14 –48.6 –44.5
FourPeople –28.8 –42.4 videoSRC15 –30.2 –28.7
Johnny –20.5 –43.3 videoSRC16 –11.8 –19.4
HEVC Class RGB EBULupoCandlelight –19.4 –28.3 videoSRC17 –29.1 –29.6
OldTownCross –33.5 –35.0 videoSRC18 12.5 16.4
ParkScene 0.3 –16.2 videoSRC19 –30.0 –6.8
DucksAndLegs –18.0 –7.9 videoSRC20 50.2 68.2
Kimono1 –29.6 –29.6 videoSRC21 –25.7 –17.1
EBURainFruits –6.0 –35.8 videoSRC22 –26.3 –24.0
UVG HoneyBee –21.7 –47.2 videoSRC23 –37.8 –31.6
ReadySteadyGo –31.2 –18.1 videoSRC24 –33.4 –30.8
Bosphorus –1.0 –29.5 videoSRC25 –21.7 –23.2
YachtRide –26.8 –26.5 videoSRC26 –26.6 –23.7
Jockey –3.0 90.9 videoSRC27 –11.4 –23.7
ShakeNDry –21.7 –24.3 videoSRC28 –20.5 –31.3
Beauty –61.3 –52.4 videoSRC29 0.2 –11.7
- - - - videoSRC30 –13.3 –35.8
Figure 17: Visualization of the reference frames and temporal contexts of DCVC-B and DCVC-DC. The first row is the $BasketballDrive\_1920\times 1080\_50$ sequence in the HEVC Class B dataset. The second row is the $RaceHorses\_832\times 480\_30$ sequence in the HEVC Class C dataset. The third row is the $BQSquare\_416\times 240\_60$ sequence in the HEVC Class D dataset. The fourth row is the $FourPeople\_1280\times 720\_60$ sequence in the HEVC Class E dataset.
V-C5 Influence of Larger GOP Size
To evaluate the capability of our DCVC-B in supporting a larger GOP size, we conduct an ablation study by setting its intra period and GOP size to 64. It is important to note that we did not fine-tune our model using longer training sequences. For comparison, we also set the intra period of DCVC-DC to 64. In Fig. 16, we compare the frame quality and bitrate of our DCVC-B scheme and those of DCVC-DC. The comparison results verify that our DCVC-B scheme can effectively support a larger GOP size and mitigate the error propagation problem.
VI Limitations and Future Work
As reported in Table II, our proposed DCVC-B exhibits different compression performance across various testing datasets. In this section, we aim to identify the limitations of DCVC-B by analyzing the characteristics of each dataset. We compare the BD-rate values of DCVC-DC and DCVC-B against the anchor HM-RA-GOP16 for each video sequence in Table VII. The comparison results reveal that DCVC-B is inferior to DCVC-DC when the testing videos have larger motion, as it struggles to obtain accurate bi-directional temporal predictions. Conversely, DCVC-B is superior to DCVC-DC when the testing videos have smaller motion, as it can leverage more useful temporal information from accurate bi-directional temporal predictions. For example, as shown in Fig. 17, when the motion between bi-directional reference frames is large, DCVC-B can only predict blurred bi-directional temporal contexts for sequences such as the $BasketballDrive\_1920\times 1080\_50$ sequence in the HEVC Class B dataset and the $RaceHorses\_832\times 480\_30$ sequence in the HEVC Class C dataset. In this case, it is difficult for DCVC-B to leverage useful temporal information from bi-directional reference frames, leading to lower compression performance. In contrast, when the motion between reference frames is small, DCVC-B can predict accurate bi-directional temporal contexts, such as the $BQSquare\_416\times 240\_60$ sequence in the HEVC Class D dataset and the $FourPeople\_1280\times 720\_60$ sequence in the HEVC Class E dataset. In this case, with more useful temporal information from bi-directional temporal contexts, DCVC-B can effectively reduce the conditional entropy, leading to higher compression performance.
Regarding the limitations of DCVC-B, we identify two critical issues that need to be addressed in the future.
  • How to obtain more accurate bi-directional temporal contexts when the motion between reference frames is large.
  • How to discriminately utilize bi-directional temporal contexts with prediction errors.
By addressing these issues, we believe that higher compression performance can be achieved for deep B-frame coding schemes.
VII Conclusion